Martha is considering a $1,000 par value bond for all the following scenarios.

C. How long would it take a bond to mature that pays a 5% annual coupon rate, has a yield to maturity of 8%, and is priced at $925?
D. What is the coupon rate for an annual coupon bond that has a yield to maturity of 8%, is priced at $845, with 13 years to maturity?
E. What is the annual coupon payment on a semiannual
Answer #1

Par value, FV = $1,000

C. How long would it take a bond to mature that pays a 5% annual coupon rate, has a yield to maturity of 8%, and is priced at $925?

FV = 1,000

PV = -925

PMT = 1,000 * 0.05 = 50

I/Y = 8

CPT N

N = 2.899435195

It would take 2.899435195 years for the bond to mature.

D. What is the coupon rate for an annual coupon bond that has a yield to maturity of 8%, is priced at $845, with 13 years to maturity?

N = 13

I/Y = 8

PV = -845

FV = 1000

CPT PMT

PMT = $60.38912019

Coupon rate = PMT/FV

Coupon rate = 60.38912019/1,000

Coupon rate = 0.06038912019

Coupon rate = 6.038912019%

Can you please upvote? Thank You :-)